### v1

@@ 4.2.1. Temporal Causal Discovery
To extract the causal relationships between historical and future exogenous variables, the Temporal Causal Discovery Module adopts a patch-wise representation strategy (Nie et al., 2023; Cirstea et al., 2022b). Specifically, each historical exogenous variable is segmented into multiple patches, and each patch is projected into a temporal token.
(1) $S_{i}=\text{PatchEmbedding}(\text{Patchify}(X_{i}^{\text{exo}},i\in\{1,...,D\})),$
@@
where $S_{i}=[{s_{i,1},s_{i,2},...,s_{i,M}}]\in\mathbb{R}^{M\times d}$ denotes the all tokens of the $i$-th channel. $M$ is the number of patches. $s_{i,j}\in\mathbb{R}^{d}$ denotes the $j$-th token of $i$-th channel. PatchEmbedding maps each patch, added by its position embedding, into a d-dimensional vector via a trainable linear projector. We then employ a standard Transformer block to model the influence weights of different patches on the future exogenous variable.
(2) $Q=S_{i}\cdot W_{q^{\prime}},\ K=S_{i}\cdot W_{k^{\prime}},\ V=S_{i}\cdot W_{v^{\prime}}$
(3) $score=\sigma(\frac{Q\cdot K^{T}}{\sqrt{d}})$
(4) $S_{i}^{\prime}=score\cdot V,$
@@
To enhance robustness, instead of directly passing the generated attention score to the Temporal Causal Injection Module, we extract and transfer the learnable parameters—specifically the query ($W_{q^{\prime}}$) and key ($W_{k^{\prime}}$) matrices—of the Multi-Head Self-Attention (MSA) mechanism that generates the attention score. These parameters serve as the temporal causality representation to be injected.
Finally, the temporal tokens $S_{i}^{\prime}$ are then processed by FeedForward layer to obtain $\hat{S}_{i}\in\mathbb{R}^{M\times d}$ and then fed into a flatten prediction head to forecast future exogenous variables $\hat{Y}^{exo}$.
(5) $\hat{Y}^{exo}_{i}=\text{Linear}(\text{Flatten}(\hat{S}_{i}))$
The prediction loss of the future exogenous variables is used as the temporal causality loss during training:
(6) $L_{t}=\left\|Y^{exo}-\hat{Y}^{exo}\right\|_{1}$
4.2.2. Temporal Causal Injection
Consistent with the approach used in the Temporal Causal Discovery Module, we first obtain patch-wise representations of the historical endogenous variables.
(7) $P_{i}=\text{PatchEmbedding}(\text{Patchify}(X_{i}^{\text{endo}},i\in\{1,...,N\})),$
@@ 4.2.2. Temporal Causal Injection
Then, we employ the Causal Transformer Block to model the influence of different endogenous variable patches on the future endogenous variable, while incorporating the causal information passed from the Temporal Causal Discovery Module. Specifically, in the attention mechanism, the input tokens are projected via original query and key matrices ($W_{q}$, $W_{k}$ and $W_{v}$) within the Causal Trmblock to obtain $Q$, $K$ and $V$, and simultaneously through ($W_{q^{\prime}}$ and $W_{k^{\prime}}$) extracted from the Temporal Causal Discovery Module to generate $Q^{\prime}$ and $K^{\prime}$.
(8) $Q^{\prime}=P_{i}\cdot W_{q^{\prime}},\ K^{\prime}=P_{i}\cdot W_{k^{\prime}}$
(9) $Q=P_{i}\cdot W_{q},\ K=P_{i}\cdot W_{k},\ V=P_{i}\cdot W_{v}$
We then compute two sets of attention scores, followed by scaling and softmax operations ($\sigma$). A learnable weighting factor $\alpha$ is then introduced to fuse the two attention scores, resulting in the final fused attention score. This mechanism explicitly injects temporal causal structure into the attention computation, enabling the model to capture causally informed dependencies among tokens along the temporal dimension.
(10) $\text{FusedScore}=\alpha\cdot\sigma(\frac{Q\cdot K^{T}}{\sqrt{d}})+(1-\alpha)\cdot\sigma(\frac{Q^{\prime}\cdot{K^{\prime}}^{T}}{\sqrt{d}})$
(11) $P_{i}^{\prime}=\sigma(\text{FusedScore})\cdot V$
@@
(12) $\alpha=\mathrm{MLP}(X^{exo}_{i})^{\top}\cdot\mathrm{MLP}(X^{endo}_{i})$
@@
(13) $\hat{Y}^{endo}_{i}=\text{Linear}(\text{Flatten}(\hat{P}_{i}))$
4.3. Channel Causality Module
Similar to the temporal dimension, the Channel Causality Module aims to model and inject causal relationships along the variable (channel) dimension. This module focuses on how historical exogenous variables influence historical endogenous variables and how such causal patterns can be transferred to enhance the prediction of future endogenous variables using future exogenous variables.
4.3.1. Channel Causal Discovery
To capture the causal effects from historical exogenous variables to historical endogenous variables, the Channel Causal Discovery Module employs a series-wise representation. Specifically, we encode each historical exogenous variables $X^{exo}_{i}$ over time into a series-wise token through a series embedding layer.
(14) $U=\text{SeriesEmbedding}(X_{i}^{\text{exo}},i\in\{1,...,D\}),$
@@ 4.3.1. Channel Causal Discovery
(15) $\mathcal{Q}=U\cdot\mathcal{W}_{q^{\prime}},\ \mathcal{K}=U\cdot\mathcal{W}_{k^{\prime}},\ \mathcal{V}=U\cdot\mathcal{W}_{v^{\prime}}$
(16) $U^{\prime}=\sigma(\frac{\mathcal{Q}\cdot\mathcal{K}^{T}}{\sqrt{d}})\cdot\mathcal{V},$
@@
(17) $)\displaystyle\hat{X}^{endo}=\text{Linear}(\text{MLP}(\hat{U}$))
Similar to Section 4.2.1, the Channel Causal Discovery Module serves two purposes: (1) it is used to define the channel causality loss, and (2) its underlying attention mechanism’s learnable parameters (query and key matrices $\mathcal{W}_{q^{\prime}}$ and $\mathcal{W}_{k^{\prime}}$) are extracted to be injected into the following channel causal injection module.
The prediction loss of the historical endogenous variables is used as the temporal causality loss during training:
(18) $L_{c}=\left\|X^{endo}-\hat{X}^{endo}\right\|_{1}$
4.3.2. Channel Causal Injection
This module uses future exogenous variables to predict future endogenous variables, while incorporating the channel-level causal representations extracted in the previous step. First, the future exogenous variables are encoded using the same series embedding strategy:
(19) $O=\text{SeriesEmbedding}(Y_{i}^{\text{exo}},i\in\{1,...,D\}),$
The Causal Trmblock is then used to model channel dependencies while injecting learned causal information.
(20) $\mathcal{Q}^{\prime}=O\cdot\mathcal{W}_{q^{\prime}},\mathcal{K}^{\prime}=O\cdot\mathcal{W}_{k^{\prime}}$
(21) $\mathcal{Q}=O\cdot\mathcal{W}_{q},\mathcal{K}=O\cdot\mathcal{W}_{k},\mathcal{V}=O\cdot\mathcal{W}_{v}$
(22) $\alpha=\text{MLP}(X^{exo})^{T}\cdot\text{MLP}(Y^{exo})$
(23) $\text{FusedScore}=\alpha\cdot\sigma(\frac{\mathcal{Q}\cdot\mathcal{K}^{T}}{\sqrt{d}})+(1-\alpha)\cdot\sigma(\frac{\mathcal{Q}^{\prime}\cdot{\mathcal{K}^{\prime}}^{T}}{\sqrt{d}})$
(24) $O^{\prime}=\sigma(\text{FusedScore})\cdot\mathcal{V}$
@@ 4.3.2. Channel Causal Injection
(25) $)\displaystyle\dot{Y}^{endo}=\text{Linear}(\text{MLP}(\hat{O}$))
Table 1. Average results on common multivariate datasets for the TSF-X task, where the inputs are ($X^{\text{endo}},X^{\text{exo}}$, and $Y^{\text{exo}}$). Red: the best, Blue: the 2nd best. Full results are provided in Table 6 of Appendix B.1.
Models DAG (ours) TimeXer TFT TiDE DUET CrossLinear Amplifier TimeKAN xPatch PatchTST DLinear
Metrics mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae
ETTh1 0.071 0.206 0.081 0.219 0.094 0.239 0.079 0.217 0.225 0.396 0.216 0.388 0.180 0.344 0.177 0.335 0.165 0.328 0.236 0.407 0.235 0.404
ETTh2 0.183 0.335 0.199 0.350 0.261 0.409 0.195 0.344 0.198 0.352 0.197 0.354 0.219 0.372 0.207 0.362 0.185 0.348 0.200 0.358 0.192 0.348
ETTm1 0.052 0.172 0.054 0.175 0.060 0.185 0.055 0.178 0.199 0.375 0.194 0.372 0.192 0.367 0.169 0.332 0.146 0.312 0.183 0.358 0.187 0.349
ETTm2 0.117 0.253 0.128 0.267 0.193 0.338 0.119 0.256 0.148 0.307 0.163 0.326 0.167 0.330 0.159 0.318 0.148 0.302 0.157 0.321 0.139 0.294
Weather 0.002 0.026 0.001 0.027 0.002 0.031 0.002 0.030 0.002 0.029 0.008 0.074 0.011 0.083 0.007 0.067 0.002 0.029 0.009 0.076 0.010 0.080
Electricity 0.194 0.329 0.208 0.337 0.375 0.471 0.313 0.417 0.281 0.411 0.264 0.399 0.277 0.413 0.301 0.429 0.265 0.400 0.269 0.404 0.288 0.420
Traffic 0.129 0.203 0.142 0.238 0.132 0.220 0.154 0.255 0.263 0.322 0.232 0.291 0.194 0.276 0.211 0.285 0.221 0.287 0.233 0.295 0.212 0.281
Exchange 0.212 0.353 0.250 0.380 0.334 0.438 0.293 0.398 0.270 0.455 0.312 0.494 0.298 0.463 0.425 0.565 0.298 0.463 0.464 0.618 0.363 0.510
4.4. Loss Function
@@
The training objective of DAG consists of three components: 1) the temporal causality loss $L_{t}$ introduced in Section 4.2.1, which measures how well the model captures temporal causal structures when forecasting future exogenous variables. 2) the channel causality loss $L_{c}$ from Section 4.3.1, which reflects the modeling error in capturing the causal relationships between historical exogenous and endogenous variables. 3) the final forecasting loss $L_{f}$ for future endogenous variables, computed based on the fused prediction, which evaluates the final prediction accuracy of the target variables.
To obtain the final prediction for future endogenous variables, we fuse two candidate predictions using a fusion weight $\lambda_{1}$:
(26) $\hat{Y}^{endo}=\lambda_{1}\cdot\ddot{Y}^{endo}+(1-\lambda_{1})\cdot\dot{Y}^{endo},$
@@
where $\ddot{Y}^{endo}$ and $\dot{Y}^{endo}$ are prediction outputs from Temporal Causal Injection Module and Channel Causal Injection Module.
The forecasting loss is defined as:
(27) $L_{f}=\left\|Y^{endo}-\hat{Y}^{endo}\right\|_{1}$
Finally, the total loss function combines the forecasting loss with the causality-related losses:
(28) $L_{total}=L_{f}+\lambda_{2}\cdot(L_{t}+L_{c}),$
@@
where $\lambda_{2}$ is a causality weight that balances the contribution of causality modeling during training.
5. Experiments
5.1. Experimental Settings
5.1.1. Datasets
To ensure comprehensive and fair comparisons across different models, 1) we perform experiments on eight common multivariate forecasting datasets (ETT with 4 subsets, Weather, Exchange, Electricity, and Traffic). Notably, although the future exogenous variables in these datasets are not known, we conduct vanilla forecasting with exogenous variables by treating the last dimension of the multivariate data as the endogenous series and the remaining dimensions as exogenous variables. 2) Additionally, we conduct experiments on 12 real-world datasets that satisfy the TSF-X conditions, including 5 EPF (Lago et al., 2021; Wang et al., 2024) datasets and 7 datasets we sourced ourselves. The future exogenous variables in these datasets are either approximately known or highly accurate. More details about the datasets are provided in Appendix A.1, Table 5.
5.1.2. Baselines
We comprehensively compare our model against 10 baselines, including 1) methods that inherently support future exogenous variables, such as TimeXer (Wang et al., 2024), TFT (Lim et al., 2021), and TiDE (Das et al., 2023), 2) as well as methods that do not originally support future exogenous variables, like DUET (Qiu et al., 2025d), CrossLinear (Zhou et al., 2025b), Amplifier (Fei et al., 2025), TimeKAN (Huang et al., 2025), xPatch (Stitsyuk and Choi, 2025), PatchTST (Nie et al., 2023), and DLinear (Zeng et al., 2023). For methods that do not support future covariates, we adapt them to incorporate future exogenous variables using an MLP fusion approach—see Algorithm 1 in Appendix A.3. Note that, for fairness, all methods in our experiments use historical endogenous variables, historical exogenous variables, and future exogenous variables.
@@ 5.1.2. Baselines
Table 2. Results on 12 real-world datasets that satisfy the TSF-X task, where the inputs are ($X^{\text{endo}},X^{\text{exo}}$, and $Y^{\text{exo}}$). Red: the best, Blue: the 2nd best. Avg means the average results from two forecasting horizons. Experimental setup is detailed in Section A.2.

### v2

@@ 4.2.1 Temporal Correlation Discovery
To extract the correlation between historical and future exogenous variables, the Temporal Correlation Discovery Module adopts a patch-wise representation strategy (Nie et al., 2023; Cirstea et al., 2022b). Specifically, each historical exogenous variable is segmented into multiple patches, and each patch is projected into a temporal token.
$S_{i}=\text{PatchEmbed}(\text{Patchify}(X_{i}^{\text{exo}},i\in\{1,...,D\})),$ (2)
@@
where $S_{i}=[{s_{i,1},s_{i,2},...,s_{i,M}}]\in\mathbb{R}^{M\times d}$ denotes the all tokens of the $i$-th channel. $M$ is the number of patches. $s_{i,j}\in\mathbb{R}^{d}$ denotes the $j$-th token of $i$-th channel. PatchEmbed maps each patch, added by its position embedding, into a d-dimensional vector via a linear projector. We then employ a standard Transformer block to model the influence weights of different patches on the future exogenous variable.
$Q=S_{i}\cdot W_{q^{\prime}},\ K=S_{i}\cdot W_{k^{\prime}},\ V=S_{i}\cdot W_{v^{\prime}},$ (3)
$score=\sigma(\frac{Q\cdot K^{T}}{\sqrt{d}}),S_{i}^{\prime}=score\cdot V,$ (4)
@@
To enhance robustness, instead of directly passing the generated attention score to the Temporal Correlation Injection Module, we extract and transfer the learnable parameters—specifically the query ($W_{q^{\prime}}$) and key ($W_{k^{\prime}}$) matrices—of the Multi-Head Self-Attention (MSA) mechanism that generates the attention score. These parameters serve as the temporal correlation representation to be injected.
Finally, the temporal tokens $S_{i}^{\prime}$ pass through a FeedForward layer to yield $\hat{S}_{i}\in\mathbb{R}^{M\times d}$, which is subsequently flattened and projected to forecast future exogenous variables $\hat{Y}^{exo}$.
$\hat{Y}^{exo}_{i}=\text{Linear}(\text{Flatten}(\hat{S}_{i})).$ (5)
The prediction loss of the future exogenous variables is used as the temporal correlation loss during training:
$L_{t}=\left\|Y^{exo}-\hat{Y}^{exo}\right\|_{1}.$ (6)
4.2.2 Temporal Correlation Injection
Consistent with the approach used in the Temporal Correlation Discovery Module, we first obtain patch-wise representations of the historical endogenous variables.
$P_{i}=\text{PatchEmbed}(\text{Patchify}(X_{i}^{\text{endo}},i\in\{1,...,N\})),$ (7)
@@ 4.2.2 Temporal Correlation Injection
Then, we employ the Correlation Transformer Block to model the influence of different endogenous variable patches on the future endogenous variable, while incorporating the correlation information passed from the Temporal Correlation Discovery Module. Specifically, in the attention mechanism, the input tokens are projected via original query and key matrices ($W_{q}$, $W_{k}$ and $W_{v}$) within the Correlation Trmblock to obtain $Q$, $K$ and $V$, and simultaneously through ($W_{q^{\prime}}$ and $W_{k^{\prime}}$) extracted from the Temporal Correlation Discovery Module to generate $Q^{\prime}$ and $K^{\prime}$.
$Q^{\prime}=P_{i}\cdot W_{q^{\prime}},\ K^{\prime}=P_{i}\cdot W_{k^{\prime}},$ (8)
$Q=P_{i}\cdot W_{q},\ K=P_{i}\cdot W_{k},\ V=P_{i}\cdot W_{v}.$ (9)
We then compute two sets of attention scores, followed by scaling and softmax operations ($\sigma$). A learnable weighting factor $\alpha$ is then introduced to fuse the two attention scores, resulting in the final fused attention score. This mechanism explicitly injects temporal correlation structure into the attention computation, enabling the model to capture Correlation informed dependencies among tokens along the temporal dimension.
$S_{\text{fused}}=\alpha\cdot\sigma(\frac{Q\cdot K^{T}}{\sqrt{d}})+(1-\alpha)\cdot\sigma(\frac{Q^{\prime}\cdot{K^{\prime}}^{T}}{\sqrt{d}}),$ (10)
$P_{i}^{\prime}=\sigma(S_{\text{fused}})\cdot V.$ (11)
@@
$\alpha=\mathrm{MLP}(X^{exo}_{i})^{\top}\cdot\mathrm{MLP}(X^{endo}_{i}).$ (12)
@@
$\hat{Y}^{endo}_{i}=\text{Linear}(\text{Flatten}(\hat{P}_{i})).$ (13)
4.3 Channel Correlation Module
Similar to the temporal dimension, the Channel Correlation Module aims to model and inject correlation along the variable (channel) dimension. This module focuses on how historical exogenous variables influence historical endogenous variables and how such correlation patterns can be transferred to enhance the prediction of future endogenous variables using future exogenous variables.
4.3.1 Channel Correlation Discovery
To capture the correlation effects from historical exogenous variables to historical endogenous variables, the Channel Correlation Discovery Module employs a series-wise representation. Specifically, we encode each historical exogenous variables $X^{exo}_{i}$ over time into a series-wise token through a series embedding layer.
$U=\text{SeriesEmbedding}(X_{i}^{\text{exo}},i\in\{1,...,D\}),$ (14)
@@ 4.3.1 Channel Correlation Discovery
$\mathcal{Q}=U\cdot\mathcal{W}_{q^{\prime}},\ \mathcal{K}=U\cdot\mathcal{W}_{k^{\prime}},\ \mathcal{V}=U\cdot\mathcal{W}_{v^{\prime}},$ (15)
$U^{\prime}=\sigma(\frac{\mathcal{Q}\cdot\mathcal{K}^{T}}{\sqrt{d}})\cdot\mathcal{V},$ (16)
@@
$).\displaystyle\hat{X}^{endo}=\text{Linear}(\text{MLP}(\hat{U}$)). (17)
Similar to Section 4.2.1, the Channel Correlation Discovery Module serves two purposes: (1) it is used to define the channel correlation loss, and (2) its underlying attention mechanism’s learnable parameters (query and key matrices $\mathcal{W}_{q^{\prime}}$ and $\mathcal{W}_{k^{\prime}}$) are extracted to be injected into the following channel correlation injection module.
The prediction loss of the historical endogenous variables is used as the temporal correlation loss during training:
$L_{c}=\left\|X^{endo}-\hat{X}^{endo}\right\|_{1}.$ (18)
4.3.2 Channel Correlation Injection
This module uses future exogenous variables to predict future endogenous variables, while incorporating the channel-level correlation representations extracted in the previous step. First, the future exogenous variables are encoded using the same series embedding strategy:
$O=\text{SeriesEmbedding}(Y_{i}^{\text{exo}},i\in\{1,...,D\}),$ (19)
The Correlation Trmblock is then used to model channel dependencies while injecting learned correlation information.
$\mathcal{Q}^{\prime}=O\cdot\mathcal{W}_{q^{\prime}},\mathcal{K}^{\prime}=O\cdot\mathcal{W}_{k^{\prime}},$ (20)
$\mathcal{Q}=O\cdot\mathcal{W}_{q},\mathcal{K}=O\cdot\mathcal{W}_{k},\mathcal{V}=O\cdot\mathcal{W}_{v},$ (21)
$\alpha=\text{MLP}(X^{exo})^{T}\cdot\text{MLP}(Y^{exo}),$ (22)
$S_{\text{fused}}=\alpha\cdot\sigma(\frac{\mathcal{Q}\cdot\mathcal{K}^{T}}{\sqrt{d}})+(1-\alpha)\cdot\sigma(\frac{\mathcal{Q}^{\prime}\cdot{\mathcal{K}^{\prime}}^{T}}{\sqrt{d}}),$ (23)
$O^{\prime}=\sigma(S_{\text{fused}})\cdot\mathcal{V}.$ (24)
@@ 4.3.2 Channel Correlation Injection
$).\displaystyle\dot{Y}^{endo}=\text{Linear}(\text{MLP}(\hat{O}$)). (25)
4.4 Loss Function
@@
The training objective of DAG consists of three components: 1) the temporal correlation loss $L_{t}$ introduced in Section 4.2.1, which measures how well the model captures temporal correlation structures when forecasting future exogenous variables. 2) the channel correlation loss $L_{c}$ from Section 4.3.1, which reflects the modeling error in capturing the correlation relationships between historical exogenous and endogenous variables. 3) the final forecasting loss $L_{f}$.
The final prediction is derived by fusing two candidates with a fusion weight $\lambda_{1}$:
$\hat{Y}^{endo}=\lambda_{1}\cdot\ddot{Y}^{endo}+(1-\lambda_{1})\cdot\dot{Y}^{endo},$ (26)
@@
where $\ddot{Y}^{endo}$ and $\dot{Y}^{endo}$ are prediction outputs from Temporal Correlation Injection Module and Channel Correlation Injection Module. The forecasting loss is defined as:
$L_{f}=\left\|Y^{endo}-\hat{Y}^{endo}\right\|_{1}.$ (27)
The total loss combines forecasting and correlation losses:
$L_{total}=L_{f}+\lambda_{2}\cdot(L_{t}+L_{c}),$ (28)
@@
where $\lambda_{2}$ is a correlation weight that balances the contribution of correlation modeling during training.
5 Experiments
In Section 5.1, we introduce the datasets, baselines, and implementation details. Section 5.2 presents the main experimental results. Section 5.3 provides a detailed model analysis, including experiments excluding future exogenous variables and ablation studies. Due to space constraints, parameter sensitivity analyses, extended lookback experiments, and visualization experiments are presented in Appendices B.2, B.3, and B.4, respectively.
@@ 5 Experiments
Table 1: Average results on 12 real-world datasets that satisfy the TSF-X task, where the inputs are ($X^{\text{endo}},X^{\text{exo}}$, and $Y^{\text{exo}}$). Red: the best, Blue: the 2nd best. Full results are available in Table 5 in Appendix B.1.